OpenStudy (anonymous):

Shanae mixes feed for various animals at the zoo so that the feed has the right amount of protein. Feed X is 18% protein and Feed Y is 10% protein. If Shanae needs a 50 lb bag that is 15% protein, how much of each feed should she use?

OpenStudy (anonymous):

Uhhhh...........

OpenStudy (anonymous):

@grray7

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@prettybunny1234567

OpenStudy (paxpolaris):

X+Y= 50 ... since the total weight is 50

OpenStudy (anonymous):

@prettybunny1234

OpenStudy (paxpolaris):

and the protein in X + the protein in Y must equal the protein in the 50 pound mix

OpenStudy (paxpolaris):

\[0.18X+0.10Y=0.15(50)\]

OpenStudy (paxpolaris):

solve the 2 equations for X and Y